A number consists of two digits.
When the number is divided by the units digit, the answer is 3.
If the digits are reversed, the new number exceeds the original number by 36.
Determine the original number.

Answer by richwmiller(17219)

x=10a+b,
x/b=3,
y=10b+a,
y-x=36
a = 1, b = 5,
x = 15, y = 51
check
15/5=3
51-15=36
ok

Just for fun how could you solve this by using logic and not algebra?
We are looking for a 2 digit product of 3 and a single digit from 0 to 9.
The factor is also the second digit
3*0=0
3*1=3
3*2=6
3*3=9
Now we have reached two digit products.
3*4=12
3*5=15
3*6=18
3*7=21
3*8=24
3*9=27
Only 3*5=15 satisfies the conditions
15 reversed =51
51-15=36
Success!